OpenStudy (anonymous):

find the critical numbers of f(x)=2xe^(6x)

OpenStudy (anonymous):

f'(c)=0,f'(x)=2(6xe^(6x)+e^(6x))=0

OpenStudy (zehanz):

I assume you mean the points where f'(x)=0: Before you can solve this equation, you must calculate the derivative of f, using the product rule:\[f'(x)=(2x)'\cdot e^{6x} + 2x \cdot (e^{6x})'\]Could you do the next step?

OpenStudy (anonymous):

x=-1/6

OpenStudy (anonymous):

thanks for your help how do i give you a medal?

OpenStudy (zehanz):

The derivative is:\[f'(x)=2e^{6x}+2x \cdot e^{6x} \cdot 6=(12x+2)e^{6x}\]Now solve f'(x)=0:\[(12x+2)e^{6x}=0 \Leftrightarrow 12x+2=0\]\[12x=-2 \Leftrightarrow x=-\frac{ 1 }{ 6 }\]

OpenStudy (zehanz):

Click "Best response" ;)

OpenStudy (anonymous):

thanks

OpenStudy (zehanz):

Thx!